### v1

# New Primitive OBZCPs of Lengths up to 49

Abstract
This paper aims to search for new primitive optimal and sub-optimal Odd Binary Z-Complimentary Pairs (OBZCPs) for lengths up to 49. As an alternative to the celebrated binary Golay complementary pairs, optimal OBZCPs are the best almost-complementary sequence pairs having odd lengths. We introduce a computer search algorithm with complexity $O(2^{N})$, where $N$ denotes the sequence length and then show optimal results for all $27\leq N\leq 33$ and $N=37,41,49$. For those sequence lengths (i.e., $N=35,39,43,45,47$) with no optimal pairs, we show OBZCPs with largest zero-correlation zone (ZCZ) widths (i.e., $Z$-optimal). Finally, based on the Pursley-Sarwate criterion, we present a table of OBZCPs with smallest demerit factors in terms of the combined auto-correlation and cross-correlation.
Index Terms:
Aperiodic correlation, Golay complementary pair (GCP), zero-correlation zone (ZCZ), Z-complementary pair (ZCP), odd-length binary ZCP (OBZCP), Pursley-Sarwate criterion.
I Introduction
I-A Background
Complementary pairs of sequences are a useful concept in coding theory, wireless communication, radar sensing, and signal processing. The general design objective is to find two equal-length sequences whose maximum aperiodic auto-correlation function (AACF) sums should be as small as possible (ideally zero). Pioneered by Marcel J. E. Golay [1] pioneered in 1951, binary complementary pair was first studied in his design of infrared multislit spectrometry, a detector which isolates the desired radiation with a fixed single wavelength from background radiation with many different wavelengths. Formally, a pair of sequences is called a Golay complementary pair (GCP) [2] if their AACF sums are zero for all the non-zero time-shifts. Since it is generally difficult to find a single unimodular sequence with zero AACF sidelobes [^1] Although a Huffman sequence has zero AACF sidelobes except for the end time-shift, its sequence elements may have different magnitudes., GCP provides a solution by allowing two sequences to work in a collaborative way. Some representative applications of the GCPs (and their extensions/variants) include: peak-to-mean envelope power ratio (PMEPR) reduction of multicarrier signals [3, 4, 5, 6], Doppler resilient radar waveforms [7, 8, 9], channel estimation [10, 11, 12, 13], intercell interference rejection [14], multicarrier code-division multiple access [15, 16, 17], etc.
Despite their wide applications, however, binary GCPs are limited to even-lengths only. More specifically, the existing binary GCPs are only known to have sequence lengths of $2^{\alpha}10^{\beta}26^{\gamma}$ (where $\alpha,\beta,\gamma$ are non-negative integers). For the odd-length case, optimal binary almost-complementary pairs are studied in [18]. Their optimization criteria is to maximize the zero-correlation zone (ZCZ) width of an odd-length binary pair, whilst minimizing its out-of-zone AACF sums. Such optimal binary pairs exhibit the correlation properties closest to GCPs and are called optimal odd-length binary Z-complementary pairs (OBZCPs). It is found that each optimal OBZCP has maximum ZCZ width of $(N+1)/2$, and minimum out-of-zone AACF sum magnitude of 2, where $N$ denotes the sequence length (odd).
By applying insertion to certain binary Golay-Davis-Jedwab (GDJ) complementary pairs [4], optimal OBZCPs [^2] In this paper, without any specific announcement, we are only interested in Type-I OBZCPs each having ZCZ around the in-phase time-shift. There are also Type-II OBZCPs where the ZCZs are centered around the end-shift positions, leading to zero correlation sidelobes away from the in-phase time-shift, but these are not our focus. of lengths $12^{m}+$ 1, where $m$ is a positive integer, are constructed in [18]. Subsequently, optimal OBZCPs of generic lengths $12^{\alpha}10^{\beta}26^{\gamma}+$ 1 (where $\alpha,~{}\beta,~{}\gamma$ are non-negative integers and $\alpha\geq 1$) are obtained in [19], thanks to certain structural properties of binary GCPs obtained from Turyn’s method [20].
I-B Contributions
Searching for new optimal OBZCPs is of interest for understanding their deeper structural properties and for providing higher level of flexibility in practical applications. Besides the aforementioned systematic constructions, it is equally important to look for short and primitive OBZCPs because they may be used to generate longer OBZCPs through certain recursive operations. For example, in 2003, Borwein and Ferguson carried out an exhaustive computer search to search new GCPs of lengths up to 100 [21]. Although some primitive OBZCPs are reported in [18], their lengths are capped to 25 only.
The main objective of this paper is thus to look for primitive optimal (or sub-optimal) OBZCPs of lengths $27\leq N\leq 49$. By leveraging the High Performance Computing (HPC) facility [^3] https://www.essex.ac.uk/staff/it-services/hpc. at the University of Essex, many new OBZCPs are found by a fine-grained computer search algorithm involving mapping, grouping, and Gray code counting. Besides the OBZCP optimality criteria on maximum ZCZ width and minimum out-of-zone AACF sums, we also use the Pursley-Sarwate criterion [22, 23] for selection of the best binary odd-length pairs with lowest combined auto-correlation and cross-correlation.
This paper is organized as follows. In Section II, we give some preliminaries and important equations/facts on OBZCPs. In Section III, we introduce the general considerations and programming techniques of the proposed algorithm. In Section IV, new OBZCPs found by computer search are presented. A list of OBZCPs with smallest demerit factors are given in Section V. Finally, this paper is concluded in Section VI.
II Preliminaries
Throughout this paper, we are interested in binary sequence pairs whose elements are drawn from the set of $\mathbb{Z}_{2}=\{0,1\}$. A length-$N$ vector is called a binary sequence if it is over $\mathbb{Z}^{N}_{2}$. For convenience, whenever necessary, binary sequences may also be shown over $\{1,-1\}^{N}$. For $\mathbf{a}=(a_{0},a_{1},\cdots,a_{N-1})$ over $\mathbb{Z}^{N}_{2}$, let ${\mathbf{a}}(z)$ be the associated polynomial of $z$ as follows,
${\mathbf{a}}(z)=\sum\limits_{\tau=0}^{N-1}(-1)^{a_{\tau}}z^{\tau}.$ (1)
For two binary sequences ${\mathbf{{a}}}$ and $\mathbf{{{{{b}}}}}$ over $\mathbb{Z}^{N}_{2}$, define
$\rho_{\mathbf{a},\mathbf{b}}(\tau)=\left\{\begin{array}[]{cl}\sum\limits_{i=0}^{N-1-\tau}(-1)^{a_{i}+b_{i+\tau}},&~{}~{}0\leq\tau\leq N-1;\\ \sum\limits_{i=0}^{N-1-\tau}(-1)^{a_{i+\tau}+b_{i}},&~{}~{}-(N-1)\leq\tau\leq-1;\\ 0,&~{}~{}|\tau|\geq N.\end{array}\right.$ (2)
When $\mathbf{a}\neq\mathbf{b}$, $\rho_{\mathbf{a},\mathbf{b}}(\tau)$ is called the aperiodic cross-correlation function of $\mathbf{a}$ and $\mathbf{b}$; otherwise, it is called the AACF. For simplicity, the AACF of a will be sometimes written as $\rho_{\mathbf{a}}(\tau)$.
II-A Binary Z-complementary pairs (ZCPs)
Definition 1
Let $\mathbf{a}$ and $\mathbf{b}$ be over $\mathbb{Z}^{N}_{2}$. $(\mathbf{a},\mathbf{b})$ is said to be a binary ZCP with ZCZ width of $Z$ if and only if [24]
$\rho_{\mathbf{a}}(\tau)+\rho_{\mathbf{b}}(\tau)=0,~{}~{}\text{for any}~{}1\leq\tau\leq Z-1.$ (3)
In this case, $\rho_{\mathbf{a}}(\tau)+\rho_{\mathbf{b}}(\tau)$ for $Z\leq\tau\leq N-1$, is called the out-of-zone aperiodic auto-correlation sum of $\mathbf{a}$ and $\mathbf{b}$ at time-shift $\tau$. When $Z=N$, a ZCP reduces to a Golay complementary pair (GCP) [2].
Fact 1
Each OBZCP $(\mathbf{a},\mathbf{b})$ has the maximum ZCZ of width $(N+1)/2$ [25], i.e., $Z\leq(N+1)/2$, where $N$ denotes the sequence length. An OBZCP is said to be $Z$-optimal if $Z=(N+1)/2$.
Fact 2
The magnitude of each out-of-zone aperiodic auto-correlation sum for a $Z$-optimal OBZCP $(\mathbf{a},\mathbf{b})$ is lower bounded by 2 [18], i.e.,
$\Bigl{|}\rho_{\mathbf{a}}(\tau)+\rho_{\mathbf{b}}(\tau)\Bigl{|}\geq 2,~{}~{}\text{for any}~{}(N+1)/2\leq\tau\leq N-1.$
A $Z$-optimal OBZCP is said to be optimal if $\Bigl{|}\rho_{\mathbf{a}}(\tau)+\rho_{\mathbf{b}}(\tau)\Bigl{|}=2$ holds for all $(N+1)/2\leq\tau\leq N-1$.
Fact 3
For an optimal (or a $Z$-optimal) OBZCP $(\mathbf{a},\mathbf{b})$, the following equations are satisfied [18]
[…]
where $21\leq r\leq(N-3)/$ 2.
II-B Pursley-Sarwate Criterion [22, 23]
For a binary sequence pair $(\mathbf{a},\mathbf{b})$ of length $N$, the cross-correlation demerit factor between $\mathbf{a}$ and $\mathbf{b}$ is defined by
$\text{CDF}(\mathbf{a},\mathbf{b})=\frac{\sum\limits_{\tau=1-N}^{N-1}|\rho_{\mathbf{a},\mathbf{b}}(\tau)|^{2}}{N^{2}}.$ (5)
For a binary sequence $\mathbf{a}$, the auto-correlation demerit factor of $\mathbf{a}$ is defined by
$\text{ADF}(\mathbf{a})=-1+\text{CDF}(\mathbf{a},\mathbf{a}).$ (6)
Furthermore, let us define the Pursley-Sarwate criterion of binary pair $(\mathbf{a},\mathbf{b})$ as follows:
$\text{PSC}(\mathbf{a},\mathbf{b})=\sqrt{\text{ADF}(\mathbf{a})\cdot\text{ADF}(\mathbf{b})}+\text{CDF}(\mathbf{a},\mathbf{b}).$ (7)
According to Pursley and Sarwate [22], we have $\text{PSC}(\mathbf{a},\mathbf{b})\geq 1$. Katz and Moore pointed out in [23] that binary GCPs of lengths $2^{\alpha}10^{\beta}26^{\gamma}$ (where $\alpha,\beta,\gamma$ are non-negative integers) satisfy the Pursley-Sarwate lower bound with equality, i.e., $\text{PSC}(\mathbf{a},\mathbf{b})=1$.
In this paper, we also use Pursley-Sarwate criterion as a sieve to select OBZCPs. Specifically, we look for optimal OBZCPs or $Z$-optimal OBZCPs with PSC values closest to 1.
III Proposed Algorithm
III-A General Considerations
In this section we describe our main considerations for our proposed algorithm with the complexity of $O(2^{N})$. Exhaustive computer search of all pairs is infeasible due to the high complexity of $O(2^{2N})$. We present a novel algorithm based on a two-step approach using the algebraic constructions and software data structures within a reasonable memory constraint.
In order to limit duplicated calculations, the first consideration is to exclude equivalent pairs. Similar to the definition in [24], two pairs $(\mathbf{a}=(a_{0},\ldots,a_{N-1}),\mathbf{b}=(b_{0},\ldots,b_{N-1}))$ are said to be equivalent, if one can be obtained by the other with one of the following operations:
  • Interchange: if $(\mathbf{a},\mathbf{b})$ is a solution, then so is $(\mathbf{b},\mathbf{a})$;
  • Negation: if $(\mathbf{a},\mathbf{b})$ is a solution, then so is $(neg(\mathbf{a}),\mathbf{b})$, where $neg(\mathbf{a})=(1-a_{0},1-a_{1},\ldots,1-a_{N-1})$;
  • Reciprocal: if $(\mathbf{a},\mathbf{b})$ is a solution, then so is $(recp(\mathbf{a}),\mathbf{b})$, where $recp(\mathbf{a})=(a_{N-1},\ldots,a_{1},a_{0})$.
Naturally, a series of these operations above lead to an equivalent solution. For instance $(neg(\mathbf{b})$, $recp(neg(\mathbf{a})))$ is also equivalent to $(\mathbf{a},\mathbf{b})$.
To proceed, without loss of generality, we assume that $a_{N-1}=b_{N-1}=1$. We then run as separate cases for the all possible combinations of the middle bits $a_{(N-1)/2}$, $b_{(N-1)/2}$ and $a_{0}=b_{0}$ due to (4).
To visualize, we consider the following OBZCP structure in our search algorithm:
$\left[\begin{matrix}a_{0}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{1}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{(N-3)/2}}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{a_{(N-1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{(N+1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{N-2}}&1\\ a_{0}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{1}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{(N-3)/2}}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{b_{(N-1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{(N+1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{N-2}}&1\end{matrix}\right].$
Next, we leverage the following three important data handling strategies in our proposed algorithm:
  • Mapping;
  • Grouping;
  • Gray counting and updates.
III-B Mapping
The purpose of mapping is to link each sequence $\mathbf{b}$ with a proper sequence $\mathbf{a}$ such that the ZCZ can be achieved. Based on equation (3), we have
$\rho_{\mathbf{b}}(\tau)=-\rho_{\mathbf{a}}(\tau)$ (8)
for each $\tau=1,\ldots,(N-3)/2$. Suppose that there are $k$ sequences of $\mathbf{a}$ which satisfy the above equation. We store all the $\rho$ values and their corresponding sequences $\mathbf{a}$ in a map $\{\rho_{\mathbf{a}}\rightarrow[\mathbf{a}_{1},\ldots,\mathbf{a}_{k}]\}$ where
$\rho_{\mathbf{a}}=(\rho_{\mathbf{a}}(1),\rho_{\mathbf{a}}(2),\ldots,\rho_{\mathbf{a}}((N-3)/2))$
for all $\mathbf{a}\in\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$ that generate the same $\rho$.
For each sequence $\mathbf{b}$, proper software implementation of such map will permit us to find all possible corresponding sequences $\mathbf{a}$ in constant $O(1)$ time.
Theoretically, with this approach we can attack the whole problem. However, this may require us to put maximally all $2^{N}$ sequences of $\mathbf{a}$ and their corresponding $\rho$ values in the memory. This is quite significant considering the physical memory limitations. Next, we introduce grouping in order to overcome this limitation.
III-C Grouping
In the first step, we group the sequence elements by using the property below:
$a_{r}+a_{N-1-r}+b_{r}+b_{N-1-r}\equiv 1~{}(\text{mod}~{}2).$ (9)
Let us define a chunk
$\mathbf{C}=(c_{1},c_{2},...,c_{(N-3)/2)})$ (10)
and we say that $\mathbf{a}\in\mathbf{C}$ iff
$c_{i}=a_{i}\oplus a_{N-1-i},i=1,\ldots,(N-3)/2$ (11)
Thus, each $\mathbf{a}$ belongs to exactly one chunk. By noting that that $a_{0},a_{(N-1)/2},a_{N-1}$ are fixed, each chunk contains $2^{(N-3)/2}$ sequences of $\mathbf{a}$ and one needs to iterate over $2^{(N-3)/2}$ chunks.
III-D Gray code counting
For every sequence $\mathbf{b}$, note that $b_{N-1}=1$, $b_{(N-1)/2}$ is fixed and $b_{0}=a_{0}$ (4). To proceed, $(b_{1},b_{2},\ldots,b_{(N-3)/2})$ is said to be the lower part of sequence $\mathbf{b}$ and $(b_{(N+1)/2},b_{(N+3)/2},\ldots,b_{(N-2)})$ its upper part. Having $2^{(N-3)/2}$ chunks, we need to iterate over the lower parts of the $2^{(N-3)/2}$ $\mathbf{b}$ sequences. Our key idea is to use a Gray code to represent the lower part, whereby the upper part is recalculated based on (4). A Gray code is characterized by having the next codeword with only one bit changed. This gives us two advantages:
  • We can update only one corresponding bit in $\mathbf{b}$’s upper part, i.e. if $b_{i}$ is changed, then $b_{N-1-i}$ is changed.
  • For such a change, the weight of sequence $\mathbf{b}$ is either increased or decreased by $2$ or stays unchanged. Hence, we can track the weight $q$ of sequence $\mathbf{b}$ without recalculating it for each modification. Having directly all possible $(p,q)$ pairs [18], we can directly exclude all the non-matching combinations.
III-E Sketch of the Proposed Algorithm
Algorithm 1 OBZCPs
1:$N\geq 5$, $N$ is odd
2:$MAX\_ACC\geq 1$, maximum correlation
3:$mid_{a,b}=(0|1)$, mid bit for $\mathbf{a}$ and $\mathbf{b}$ correspondingly
4:$lsb=(0|1)$, least significant bit for $\mathbf{a}$ and $\mathbf{b}$
5:$CHUNK\_SIZE\Leftarrow(N-3)/2$
6:$NUMBER\_OF\_CHUNKS\Leftarrow 1<<CHUNK\_SIZE$
7:for $\mathbf{c}\leftarrow 1$ to $NUMBER\_OF\_CHUNKS$ do
8:     for  $\mathbf{g}\leftarrow 0$ to $12^{CHUNK\_SIZE}-$ 1 do
9:         $\mathbf{a}\leftarrow(lsb,\mathbf{g},mid_{a},rev(\mathbf{g},\mathbf{c}),1)$
10:         Add $\mathbf{a}$ to $map[\rho(\mathbf{a})]$
11:     end for
12:     $\mathbf{b\_lower}\leftarrow(1,0,0,\ldots,0)$
13:     Calculate $\mathbf{b\_upper}$ such that Eq.(4) holds for $\mathbf{b}=(lsb,\mathbf{b\_lower},mid_{b},\mathbf{b\_upper},1)$
14:     for $j\leftarrow 1$ to $2^{CHUNK\_SIZE}$ do
15:         $\rho_{\mathbf{a}}(\tau)\leftarrow-\rho_{\mathbf{b}}(\tau),\tau=1,\ldots,(N-3)/2$
16:         for $\mathbf{a}\in map[\rho(\mathbf{a})]$ do
17:              if  $\mathbf{a}\geq\mathbf{b}$ then
18:                  Print $(\mathbf{a},\mathbf{b})$ if out-of-zone correlation $\leq MAX\_ACC$
19:              end if
20:         end for
21:         Generate next Gray code for $\mathbf{b\_lower}$ part by changing single bit $x$
22:         Update $\mathbf{b\_upper}$ part by changing bit $N-1-x$
23:     end for
24:end for
25:Filter out equivalent pairs
The proposed algorithm is based on one outer loop to iterate over all the chunks and two consecutive inner loops for 1) loading data in memory map and 2) iterating over Gray code by changing a single bit on each step and retrieving matching values from the map such that equation (4) holds.
Considering that the corresponding search in table between $\mathbf{a}$ and $\mathbf{b}$ has lookup complexity of $O(1)$, iterating over $2^{(N-3)/2}$ values and $2^{(N-3)/2}$ chunks results in total complexity of $O(2^{N})$. We execute this algorithm $8$ times with all possible combinations of $a_{(N-1)/2},b_{(N-1)/2},a_{0}$.
In the algorithm, $rev(\mathbf{g},\mathbf{c})$ is defined such that $\mathbf{g}\oplus rev(\mathbf{g},\mathbf{C})\in\mathbf{C}$.
Note that the most inner loop is usually of size $0,1,2,$ rarely up to $4$ and hence does not impact the overall complexity.
IV Computer Search Results
In this section, we present optimal OBZCPs from computer search, and in case they do not exist, $Z$-optimal OBZCPs for $27\leq N\leq 49$, in Tables I to XII.
Optimal OBZCPs up to $N=25$ are reported in [18] and some optimal pairs for $N=33,41$ can be obtained via the constructions in [19]. During this computer search, we have found only two non-equivalent optimal pairs for $N=37$. We have also found optimal pairs of length $N=49$, but the search is non-exhaustive as we reached the limit of our computational resources.
Our computer search shows that there are no optimal OBZCPs for $N=35,39,43,45,47$ and this motivates us to search $Z$-optimal pairs with the maximum out-of-zone aperiodic autocorrelation sum having magnitude of $6$. In this work, we only present a few $Z$-optimal pairs in the tables below for $N=35,39,43,45,47$.
It is noted that all the OBZCPs are presented in hexadecimal notation. This allows us to map every hexadecimal digit to four binary digits, i.e. ($000\rightarrow 0$ $0$, $1\rightarrow 0001$, …, $F\rightarrow 1111$) and by stripping initial zeros if necessary. For instance $24159FE$ 24 represents $1010110011111111000100100$.
TABLE I: Optimal OBZCPs for $N=27$
Column 1 Column 2 Column 3 Column 4
(6AC2984, 42265F0) (419B094, 4038DAA) (5CBF6C2, 48798A8) (74E2BC8, 5FD9A5C)
(72C2320, 6581DAA) (57374F2, 4F2B700) (5CB287E, 409159C) (6A5F3C8, 49811D4)
(6A92700, 41D3994) (789A684, 4774284) (6DDE82E, 5230C2E) (5286774, 406832E)
(77724F1, 71B4ABF) (623AFC9, 4C14A0D) (668A84F, 4F0B77B)
TABLE II: Optimal OBZCPs for $N=29$
Column 1 Column 2 Column 3 Column 4
(144E4E10, 114693FA) (1EE4E4BA, 1BEC3950) (17CA7B3A, 14640A7C) (1FCADB38, 1C64AA7E)
(1D60D190, 129F5190) (15B04C28, 153183F6) (1F1AE682, 18AD64C0) (1A9FF192, 16CE3F2A)
(1835D190, 17CA5190) (1A9FDB38, 15605B38) (1F6CC570, 11D1AD20) (15C66FDA, 15C3E424)
(18D7E164, 1560232C) (1FCAF192, 16CE2A7E) (1D60FB3A, 146420D6) (13CDD580, 1185A836)
(1AEF4E13, 16F06EEB) (1045E4B9, 10446B47)
TABLE III: Optimal OBZCPs for $N=31$
Column 1 Column 2 Column 3 Column 4
(4ED3AE80, 43856426) (692FCE8C, 6479042A) (6EF9C5A5, 5D3703BD) (6C7FC945, 5EB32F1D)
(64AFE6B9, 640ACBC7) (779DA917, 44536F0F) (741985B7, 46D563EF) (76D0CC17, 46D563EF)
(5B4302B9, 4E054C13)
TABLE IV: Optimal OBZCPs for $N=33$
Column 1 Column 2 Column 3 Column 4
(1EA6D8C0A, 1EA6C73F4) (140C726A0, 140C6D95E) (1935FF538, 1935E0AC6) (139F55F92, 139F4A06C)
(180CBE6AC, 180CA1952) (12A614C06, 12A60B3F8) (180B39EAC, 180B26152) (12A193406, 12A18CBF8)
(18CBFEAD8, 18CBE1526) (126154072, 12614BF8C) (18CBFEAD8, 136AF059C) (1F8B39ED4, 1F8B2612A)
(15219347E, 15218CB80) (1F8CBE6D4, 1F8CA192A) (152614C7E, 15260B380) (1C377DD68, 1C3762296)
(169DD77C2, 169DC883C) (1828D8272, 17D728272) (1DC97C288, 1DC963D76) (1763D6822, 1763C97DC)
(1E72A06CA, 1A6C05630)
TABLE V: $Z$-optimal OBZCPs for $N=35$
Column 1 Column 2 Column 3
(7905A9444, 710C1A3B2) (72E2F6394, 5DE937410) (4AC870914, 40DDE22C2)
(54C870928, 40DDE22C2) (793B3EE96, 5C6F5043A) (44EA20C2D, 42D04DDE3)
Since $N=35$ is the first odd length that does not have optimal pairs, we present on the left-hand-side of Fig. 1 on the magnitude plot of the AACF sums of the pair $(7905A9444,710C1A3B2)$. On the right-hand-side of Fig. 1, we show the magnitude plot of the optimal pair $(15BCD1FAF3340,\\ 10599EA0E984A)$ with length 49.
TABLE VI: Optimal OBZCPs for $N=37$
Column 1 Column 2
(1D29F4D110, 11273940E8) (17B506C9C4, 144430A7C2)
TABLE VII: $Z$-optimal OBZCPs for $N=39$
Column 1 Column 2 Column 3
(69B1294470, 5AEF8C06E8) (6CC43E1164, 4A75EC2828) (70122A279C, 5B235A9E08)
(72577E7298, 6E0592287A) (59F26A2072, 44AF3882D0) (78C84D1254, 77E235C7D2)
TABLE VIII: Optimal OBZCPs for $N=41$
Column 1 Column 2 Column 3
(1A2903A133C, 15D6FCA133C) (1BE159959F6, 1BE1586A608) (114BF33F35C, 114BF2C0CA2)
(18215995906, 1821586A6F8) (128BF33F3AC, 128BF2C0C52) (1E08CBE6EB4, 1E08CA1914A)
(14A2614C41E, 14A260B3BE0) (1808D386EAC, 1808D279152) (12A2792C406, 12A278D3BF8)
(18A5F99F9AE, 175A069F9AE) (1945F99F9D6, 16BA069F9D6) (1F5EC9C62FA, 10A136C62FA)
(12B821E8CD2, 12B8201732C) (18128B42678, 18128ABD986) (1CEBE9CE816, 131416CE816)
(18504D857B2, 18504C7A84C) (12FAE72FD18, 12FAE6D02E6) (1A5DCFA6EF0, 15A230A6EF0)
(195DC396EFC, 16A23C96EFC) (13F7693C456, 12BB872DDF8) (1F51C9C6DFA, 10AE36C6DFA)
(1A2903C337A, 15D6FCC337A) (1E14DDB4188, 11EB22B4188) (1C9EF59CBA0, 13610A9CBA0)
(16345F3610A, 15EF27F458C) (1A774FA7BB0, 1588B0A7BB0) (19774397BBC, 1688BC97BBC)
(19FA296B9F9, 1605D66B9F9) (13DDE93D116, 12EE872F778)
TABLE IX: $Z$-optimal OBZCPs for $N=43$
Column 1 Column 2 Column 3
(667A828B448,41FB9CB0B44) (5C43DF5FA66,443D3C9BA94) (6B171FE57DC,596866ED704)
(48A06A9213E,4520EE65E18) (5862CFF91DA,557ED6C727C) (7891AACF658,4A724E0EA00)
TABLE X: $Z$-optimal OBZCPs for $N=45$
Column 1 Column 2 Column 3
(1C3009533530,15C8AF20693C) (169AA3F99F9A,12C795CBF720) (16B772EBE890,1675B3847F0E)
(1ADB163A02A9,118C8F16C00D) (13589BC02838,11EF69162A6E) (1EDDBC719A8A,16807CAE1236)
TABLE XI: $Z$-optimal OBZCPs for $N=47$
Column 1 Column 2 Column 3
(697954B3441C,4F4FEE620DD0) (776B6C9CE0AC,5447BF068530) (73686A8106F6,6145C274232C)
(5D9F943CC229,4BBA9D0B6FE3) (5375CCE48E1E,4752A49003F4) (7815ACEB273E,579B7779603A)
TABLE XII: Optimal OBZCPs for $N=49$
Column 1 Column 2
(15BCD1FAF3340,10599EA0E984A) (1F167B50599EA,1AF3340A432E0)
(1FB67150F994A,1A533E0AE3240) (151CDBFA533E0,151CDBF5ACC1E)
(1564E7FF86694,152CC3E031B2A) (156CC7FF8E494,1524E3E03992A)
(1FC66D5524E3E,1F8E494A93380) (1FCE4D552CC3E,1F86694A9B180)
(15BCD1FFA6614,150CCBE0E984A) (1FB67155ACC1E,1F066B4AE3240)
(151CDBFF066B4,151CDBF0F994A) (1F167B5AF3340,10599EAA432E0)
(1564E7FAD33C0,107996A031B2A) (156CC7FADB1C0,1071B6A03992A)
(1FC66D5071B6A,1ADB1C0A93380) (1FCE4D507996A,1AD33C0A9B180)
Figure 1: Plots of AACF sums for OBZCPs, i.e., $(7905A9444,710C1A3B2)$ and $(15BCD1FAF3340,10599EA0E984A)$, of lengths 35 and 49, respectively.
V OBZCPs with Smallest Demerit factors
Finally, we leverage the Pursley-Sarwate criterion and present OBZCPs with smallest demerit factors for sequence lengths up to 49 in Table XIII.
TABLE XIII: OBZCPs with smallest demerit factors
N PSC($g_{1}$, $g_{2}$) ADF($g_{1}$) ADF($g_{2}$) CDF($g_{1}$, $g_{2}$) $g_{1}$ $g_{2}$
3 1.4444 1.1111 1.1111 0.33333 7 5
5 1.2997 0.48 0.8 0.68 1E 16
7 1.185 0.77551 0.28571 0.71429 5D 4F
9 1.1636 0.79012 0.39506 0.60494 1E8 14C
11 1.149 0.41322 0.67769 0.61983 7C6 5DA
13 1.1005 0.071006 0.30769 0.95266 1F35 1709
15 1.1107 0.38222 0.20444 0.83111 612E 4C0A
17 1.0983 0.38754 0.609 0.61246 1FCA9 1A6E3
19 1.0931 0.40443 0.27147 0.76177 7A284 78CDA
21 1.0819 0.29932 0.46259 0.70975 1B6A87 109883
23 1.0762 0.41966 0.58601 0.58034 78B519 76D1DF
25 1.0740 0.4224 0.5248 0.6032 159FB70 11DA0CA
27 1.0664 0.36488 0.25514 0.76132 5CB287E 409159C
29 1.0608 0.27111 0.39477 0.73365 144E4E10 114693FA
31 1.0571 0.4308 0.30593 0.69407 64AFE6B9 640ACBC7
33 1.0554 0.33792 0.24977 0.76492 1E72A06CA 1A6C05630
35 1.0792 0.19102 0.25633 0.85796 44EA20C2D 42D04DDE3
37 1.0515 0.30095 0.35354 0.72535 1D29F4D110 11273940E8
39 1.0874 0.27745 0.3879 0.75937 78C84D1254 77E235C7D2
41 1.0460 0.38548 0.31886 0.69542 19FA296B9F9 1605D66B9F9
43 1.0964 0.38615 0.33423 0.73716 668D847F75A 62EEF0F6BB4
45 1.0748 0.37728 0.34963 0.7116 1ADB163A02A9 118C8F16C00D
47 1.0972 0.32866 0.41195 0.72929 5D9F943CC229 4BBA9D0B6FE3
49 1.0388 0.33986 0.39983 0.67014 156CC7FF8E494 1524E3E03992A
One can see that the PSC value of each pair is close to $1$, indicating that each pair is almost complementary. We also notice a trend that the PSC values for optimal OBZCPs generally decrease for larger sequence lengths with just one exception for $N=13$.
VI Conclusions
This paper has introduced a number of new primitive optimal and $Z$-optimal OBZCPs for $27\leq N\leq 49$ which are obtained through a computer search. By using selected algebraic properties of OBZCPs and with certain programming techniques, we have presented an algorithm with complexity of $O(2^{N})$. Next to the known OBZCP constructions for $N=33$ and $N=41$, we have found optimal sequence pairs for $N=37$ and $N=49$. Additionally, a list of OBZCPs with smallest demerit factors have been reported.
An interesting future work of this research is to develop systematic optimal (or sub-optimal) OBZCP constructions by exploiting the obtained short sequence pairs. Another ambitious future direction is to carry out an exhaustive computer search for OBZCPs up to length 100 and analyze their deeper structural properties.